During its first year of operations, the McCormick Company incurred the following manufacturing costs: Direct materials, $5 per unit, Direct labor, $2 per unit, Variable overhead, $4 per unit, and Fixed overhead, $390,000. The company produced 39,000 units, and sold 30,000 units, leaving 9,000 units in inventory at year-end. Income calculated under variable costing is determined to be $415,000. How much income is reported under absorption costing?

Multiple Choice

  • $415,000

  • $325,000

  • $805,000

  • $505,000
Answer #1

Fixed Overhead per Unit= $390000/39000=$10

Inventory in Hand :- 9000

Fixed Cost applied on Inventory : (9000*10)= $90000

Income as per Absorption costing = Income from Variable costing + Fixed COst applied on Inventory

'= $415000+90000=$505000

Hence, $505000 is correct answer.

Dear student kindly let me know, if you are having any doubt and please mark with positive rating.